Test listcase 1 — succeeded. Test dolist 1 — succeeded. Test ReverseList 1 — succeeded. Test ReverseList 2 — succeeded. Test ReverseList 1a — succeeded. Test ReverseList 2a — succeeded. Test InsertThousandsMarks 1 — succeeded. Test listcase 2 — succeeded. Test listcase 3 — succeeded. Test ziplist 1 — succeeded. Test dolist 2 — succeeded. Test dolist 3 — succeeded. Test dolist ifXXX,else,fi — succeeded.